# Latent Diffusion Model-Enabled Real-Time Semantic Communication Considering Semantic Ambiguities and Channel Noises

### v1

@@ III Deep JSCC for unknown errors and distributions
In this section, the selection of the encoder and decoder for the proposed robust and high-quality JSCC is further detailed. In Subsection III-A, WGAN and its higher quality variants are introduced to serve as the decoders for receivers. Subsection III-B firstly presents the joint training algorithm of VAE and WGAN, and Subsection III-C then provides an improved encoder that is robust to measuring errors. The fast and reliable SemCom approach for data of unknown distribution is implemented in Subsection III-D by out-of-distribution latent space exploration.
III-A Decoder: Wassertein GAN and other variants
In recent years, significant progresses have been made in the generation of ultra-high quality texts, images, and videos in the field of AIGC. The DMs, as an iconic generative model in AIGC, has also received widespread attention. Nevertheless, considering the basic characteristics of the original DMs, SemCom systems built directly based on DMs cannot effectively improve communication effciency, i.e., reduce the bandwidth requirements. Therefore, GANs with slightly lower generation quality than DMs are still selected as the decoder for the proposed JSCC. GAN is formulated based on zero-sum game between a discriminator $D_{\bm{\gamma}}(.)$ and a generator $G_{\bm{\psi}}(.)$ with the adversarial training objective
$\min_{\bm{\psi}}\max_{\bm{\gamma}}$ $\mathbb{E}_{\bm{x}\sim q(\bm{x})}\left[\log D_{\bm{\gamma}}(\bm{x})\right]$ (2)
$+\mathbb{E}_{\bm{z}\sim q_{\bm{\psi}(\bm{z})}}\left[\log\left(1-D_{\bm{\gamma}}(G_{\bm{\psi}}(\bm{z}))\right)\right],$
where $q_{\bm{\psi}}(\bm{z})$ is the prior distribution of latent vector $\bm{z}$ and $q_{\bm{\psi}}(\bm{z})=\mathcal{N}(\bm{0},\bm{I})$ in GANs. Unfortunately, although GAN can generate higher quality data compared to other generative models such as VAE, it still faces some issues such as training instability and collapse mode. To overcome these challenges, WGAN is established by replacing Kullback-Leibler divergence $\mathcal{D}_{KL}$ and Jensen-Shannon divergence $\mathcal{D}_{JS}$ with Wassertein distance $\mathcal{D}_{\mathcal{W}}$. In this way, the optimization objective of WGAN can be decoupled as
$\min_{\bm{\psi}}\mathbb{E}_{\bm{z}\sim q_{\bm{\psi}}(\bm{z})}\left[\mathcal{D}_{\mathcal{W}}\left(\bm{x}\parallel G_{\bm{\psi}}(\bm{z})\right)\right]\Leftrightarrow-\mathbb{E}_{\bm{z}\sim q_{\bm{\psi}}(\bm{z})}\left[D_{\bm{\gamma}}\left(G_{\bm{\psi}}(\bm{z})\right)\right]$ (3)
$\min_{\bm{\gamma}}-\mathbb{E}_{\bm{x}\sim q(\bm{x})}\left[D_{\bm{\gamma}}(\bm{x})\right]+\mathbb{E}_{\bm{z}\sim q_{\bm{\psi}}(\bm{z})}\left[D_{\bm{\gamma}}\left(G_{\bm{\psi}}(\bm{z})\right)\right].$
@@ III-A Decoder: Wassertein GAN and other variants
Furthermore, the variants of GAN that have been proposed for better perceptual reconstruction can also be utilized as the JSCC decoder. Among them, StyleGAN [50], Diff-GAN trained based on DMs [51], and Diff-Instuct GAN [52] distillated from DMs have achieved impressive generation resluts, and the latter two models are even comparable to the DMs in some datasets. Consequently, when receivers have higher requirements on decoding data quality, WGAN can be flexibly replaced with these superior variants, as they have the same architecture as the original GANs.
III-B Encoder: GAN Inversion by VAE-WGAN
@@
Compared to VAE, GANs are skilled in generating data with high-resolution. Nonetheless, the task of SemCom is to ensure that the signals received by receivers can accurately convey the meaning, while minimizing the bandwitdth of SemCom. For this reason, JSCC requires an encoder to determine the latent bottlenecks of the transmitted data, also known as GAN inversion [53]. Commonly, the solution of GAN inversion is to utilize a neural network-based encoder to find the optimal latent vector $\bm{z}$ given the transmitted data $\bm{x}$. Ignoring the channel’s cross entropy term of the latent space and taking into account the receiver reconstruction term and the transmitter encoding entropy, the variational upper bound can be transformed into
@@
where the proof can be seen in Appendix A. Apparently, the term $\mathbb{E}_{q_{\bm{\psi}}(\bm{z})}\left[-\log p_{\bm{\psi}}(\bm{x})\right]$ in Eq. (4) can be replaced with the training objective of the generator $G_{\bm{\psi}}(.)$ of WGAN, and term $\mathbb{E}_{q(\bm{x})}\left[\mathcal{D}_{KL}(q_{\bm{\phi}}(\bm{z}|\bm{x})\parallel p_{\bm{\psi}}(\bm{z}|\bm{x}))\right]$ indicates that the encoding latent vector $\bm{z}$ should be as consistent as possible with the input latent space of generator $G_{\bm{\psi}}(.)$ under the same transmitted data $\bm{x}$. Like VAE, the output of encoder can be represented as $q_{\bm{\phi}}(\bm{z}|\bm{x})\sim\mathcal{N}(\bm{\mu},\bm{\sigma}^{2})$, and $\bm{z}$ is reparameterized as $\bm{z}=\bm{\mu}+\bm{\sigma}\odot\bm{\epsilon}$, where $\bm{\epsilon}\sim\mathcal{N}(\bm{0},\bm{I})$ and $\odot$ denotes the element-wise product. Consequently, by combining the decoupled optimization objectives of WGAN and Eq. (4), the optimization objectives for the encoder, generator, and discriminator of JSCC are
  • $\bullet$
    Encoder $E_{\bm{\phi}}(.)$: $\mathbb{E}_{q}\big{[}\alpha_{\bm{\phi}}\mathcal{D}_{KL}\big{(}E_{\bm{\phi}}(\bm{x})\sim\mathcal{N}(\bm{\mu},\bm{\sigma}^{2})\parallel\bm{z}\sim\mathcal{N}(\bm{0},\bm{I})\big{)}+(1-\alpha_{\bm{\phi}})\mathcal{D}_{KL}\left(G_{\bm{\psi}}(E_{\bm{\phi}}(\bm{x}))\parallel\bm{x}\right)\big{]}$,
  • $\bullet$
    Generator $G_{\bm{\psi}}(.)$: $\mathbb{E}_{q}\big{[}\alpha_{\bm{\phi}}\mathcal{D}_{KL}\left(G_{\bm{\psi}}(E_{\bm{\phi}}(\bm{x}))\parallel\bm{x}\right)+(1-\alpha_{\bm{\psi}})\left(-D_{\bm{\gamma}}(G_{\bm{\psi}}(\bm{z}))-D_{\bm{\gamma}}(G_{\bm{\psi}}(E_{\bm{\phi}}(\bm{x})))\right)\big{]}$,
  • $\bullet$
    Discriminator $D_{\bm{\gamma}}(.)$: $\mathbb{E}_{q}\big{[}-D_{\bm{\gamma}(\bm{x})}+D_{\bm{\gamma}}(G_{\bm{\psi}}(\bm{z}))+D_{\bm{\gamma}}(G_{\bm{\psi}}(E_{\bm{\phi}}(\bm{x})))\big{]}$,
where $\alpha_{\bm{\phi}}$ and $\alpha_{\bm{\psi}}$ are the loss balance hyperparameters. The training process of deep CNN based VAE-WGAN with gradient penalty [54] is illustrated in Appendix B.
III-C Robust Encoder
As discussed in Section II, VAE-WGAN based SemCom systems suffer from inevitable vulnerabilities, i.e., the covert errors of sampled or measured data $\bm{x}$ or latent vectors $\bm{z}$ transmitted in wireless channels can lead to the receiver’s decoding outputs $\hat{\bm{x}}$ that are inconsistent with the original data $\bm{x}$, ultimately resulting in semantic errors. The adversarial attack methods explores the vulnerabilities of neural networks utilized for classification and regression tasks, with the goal of determining a sufficiently small and unnoticed measurement error $\bm{\delta}$ to mislead the classification or regression results. The unified optimization objective of adversarial attacks is given by
@@ III-C Robust Encoder
s.t.: $\bm{f}(\bm{x}+\bm{\delta})=target,\quad\bm{L}\leq\bm{x}+\bm{\delta}\leq\bm{U},$
where $\bm{d}(.,.)$ is the distance function, $\bm{f}(.)$ is the attacked nerual network, target denotes the output of the DL model, $\bm{L}$ and $\bm{U}$ represent the physical lower and upper bounds of input data $\bm{x}$, respectively. Specifically, in classification tasks, $target$ is the class that is inconsistent with original category of $\bm{x}$, e.g., the hackers can exploit objective (5) to make their cyber attacks undetectable by network $\bm{f}(.)$. In regression tasks, $target$ represents the output data that is not the same as the original regression results, e.g., when transmitting a digital image in SemCom systems, the receiver’s decoder may reconstruct another type of digital image, resulting in semantic ambiguity and seriously reducing the reliability of security of SemCom systems.
In order to address the challenges of semantic errors, SemCom systems should have a robust and enhanced encoder that can handle those outliers. The objectives for the sufficiently small semantic error that leads to the maximum encoding error or reconstruction error are given by
$\max_{\bm{\delta}}$ $\bm{d}\left(E_{\bm{\phi}}(\bm{x}),E_{\bm{\phi}}(\bm{x}+\bm{\delta})\right)$ (6)
s.t.: $E_{\bm{\phi}}(\bm{x}+\bm{\delta})\sim\mathcal{N}(\bm{0},\bm{I}),\left\|\bm{\delta}\right\|_{p}\leq\varepsilon,\bm{L}\leq\bm{x}+\bm{\delta}\leq\bm{U}$
and
respectively, where .$\left\|.\right\|_{p}$ denotes the p-norm. In this way, objective (6) utilizes the vulnerability of the encoder $E_{\bm{\phi}}(.)$, while objective (7) simultaneously utilizes the vulnerabilities of both the encoder $E_{\bm{\phi}}(.)$ and generator $G_{\bm{\psi}}(.)$. When solving Eq. (7), its objective can be transformed into a standard convex optimization problem
where $\lambda$ is the penalty coefficient. The constrainted convex optimization problem can be solved by using the projected gradient descent (PGD) [55] iterative optimization method to obtain semantic error $\bm{\delta}$. Consequently, the semantic error at the i-th iteration $\bm{\delta}^{i}$ is denoted by
where $P_{C}(\bm{\varsigma}^{i})$ represents the projection of $\bm{e}(\bm{\delta})$ on the set of constraints $C$, i.e., $\bm{\delta}^{i}=P_{C}(\bm{\varsigma}^{i}):={\arg\min}_{\bm{\delta}\in C}\frac{1}{2}\left\|\bm{\delta}-\bm{\varsigma}^{i}\right\|_{2}^{2}$. Let $\bm{z}^{\prime}$ be the erroneous latent vector encoded from data containing semantic errors $\bm{x}^{\prime}=\bm{x}+\bm{\delta}$, the robust variational upper bound for semantic errors is defined as
$\mathbb{E}_{q}\left[-\log p_{\bm{\psi}}(\bm{x})\right]=\mathbb{E}_{q}\left[-\log p_{\bm{\psi}}(\bm{x},\bm{x}+\bm{\delta})d(\bm{x}+\bm{\delta})\right]$ (10)
$+\mathbb{E}_{q(\bm{z}^{\prime})}\left[-\log p_{\bm{\psi}}(\bm{z}^{\prime})\right]+\frac{\beta}{2}\mathbb{E}_{q(\bm{z},\bm{z}^{\prime})}\bm{d}(\bm{z},\bm{z}^{\prime})$
$-\mathbb{E}_{q(\bm{z},\bm{z}^{\prime})}\left[-\log q(\bm{z},\bm{z}^{\prime})\right],$
where $\beta$ denotes the nonnegative coupling parameter. Evidently, the first and second term of Eq. (10) have been addressed in VAE-WGAN based JSCC, and the third term has also been optimized by sloving the semantic errors $\bm{\delta}$. For this reason, the training objective of robust encoder is
$\min_{\bm{\phi}^{\prime}}\quad\frac{\beta}{2}\mathbb{E}_{q(\bm{z},\bm{z}^{\prime})}\bm{d}(\bm{z},\bm{z}^{\prime})+\mathbb{E}_{q(\bm{z},\bm{z}^{\prime})}\left[\log q(\bm{z},\bm{z}^{\prime})\right].$ (11)
In [56], the objective (11) is equivalent to minimizing the Wassertein distance between $\bm{z}$ and the incorrect $\bm{z}^{\prime}$. Nevertheless, the ultimate goal of SemCom is to accurately reconstruct the transmitted data. Therefore, as the parameter $\bm{\psi}$ is fixed, the optimal robust encoder parameter $\bm{\phi}^{\prime}$ considering both encoder and decoder vulnerabilities is
Input: Dataset $q(\bm{x})$, learning rate $\eta_{1}$ and $\eta_{2}$, original encoder $E_{\bm{\phi}}(.)$, generator $G_{\bm{\psi}}(.)$,
Output: The updated robust encoder $E_{\bm{\phi}^{\prime}}(.)$
@@ 10      Determine $\bm{\delta}$ by $\bm{\delta}\leftarrow\bm{\delta}^{k}$;
Return Robust GAN inversion $E_{\bm{\phi}^{\prime}}(.)$
Algorithm 1 Training algorithm of robust GAN inversion $E_{\bm{\phi}^{\prime}}(.)$
III-D Out-of-Distribution Latent Space
@@
Currently, one of the biggest application bottlenecks of DL-based SemCom is their generalization performance, which means that these machine learning models will significantly degrade their performances when facing data types that are not included in the training dataset. Specifically, the data type within the training dataset is referred to as in domain, while data with an unknown distribution for those models can be referred to as out-of-domain. For the proposed VAE-WGAN architecture based SemCom, when the transmitter wants to send an out-of-distribution data, robust encoder $E_{\bm{\phi}^{\prime}}$ may encode an abnormal latent vector, and the decoder at the receiver will reconstruct data that is semantically different from the transmitted data. For this reason, when facing these unknown distributions, JSCC should improve its generalization abilities and be able to quickly adapt to search for the optimal out-of-distribution latent space.
In addition to the learning-based approaches, another approach of GAN inversion can be established by convex optimization. Given the generator $G_{\bm{\psi}}(.)$ pre-trained in a specific domain and a data $\bm{x}^{\prime\prime}$ following an unknown distribution, the optimization objective [57] for the optimal latent vector corresponding to $G_{\bm{\psi}}(.)$ is denoted by
$\min_{\hat{\bm{z}}}$ $\bm{d}\left(\bm{x}^{\prime\prime},G_{\bm{\psi}}(\hat{\bm{z}})\right)+\rho(\hat{\bm{z}})$ (13)
s.t.: $\hat{\bm{z}}\sim\mathcal{N}(\bm{0},\bm{I}),$
@@
where $\rho(.)$ is a penalty term aimed at making $\hat{\bm{z}}$ follow the prior distribution of generator $G_{\bm{\psi}}(.)$, i.e., $\hat{\bm{z}}\sim q_{\bm{\psi}}(\bm{z})$. Similar to the convex optimization problem of semantic error $\bm{\delta}$, objective (13) can also be solved through PGD without using an additional encoder. Nonetheless, this convex optimization problem can take tens to hundreds of iterations each time, so it cannot guarantee the timeliness of SemCom. To address this issue, a learning-based adaptor constructed by a lightweight single-layer neural network is utilized for out-of-distribution latent space determination. Considering the characteristics of VAE-WGAN, as shown in Fig. 3, the adapter $g_{\bm{\omega}}(.)$ parameterized by $\bm{\omega}$ is placed between the robust encoder and generator. Subsequently, when the transmitted data follows an unknown distribution, adaptor $g_{\bm{\omega}}(.)$ can perform one-shot learning to transform the latent vector $\bm{z}$ encoded by $E_{\bm{\phi}^{\prime}}(.)$ into
@@
where $\bm{b}$ denotes the bias of adaptor $g_{\bm{\omega}}(.)$. In order to improve the data quality of reconstruction, inspired by the adversarial training startegy of WGAN, this paper considers another adaptor $d_{\bm{\nu}}(.)$ composed of a lightweight fully connected (FC) layer for adversarial training with $g_{\bm{\omega}}(.)$. As illustrated in Fig. 3, during online training, the FC layer of discriminator $D_{\bm{\gamma}}(.)$ is replaced by $d_{\bm{\nu}}(.)$, and the original discriminator after removing the FC layer is denoted by $d_{\bm{\gamma}}(.)$. In this way, the online training process of $g_{\bm{\omega}}(.)$ is similar to WGAN’s training approach as illustrated in Algorithm 2.
@@
Input: Data following unknown distribution $q(\bm{x}^{\prime\prime})$, learning rate $\eta$, gradient penalty coefficient $\lambda$, robust encoder $E_{\bm{\phi}^{\prime}}(.)$, generator $G_{\bm{\psi}}(.)$, discriminator $D_{\bm{\gamma}}(.)$ and the parameters of last layer (L-th layer) is denoted by $\bm{\gamma}^{L}$
Output: The online-updated adaptor $g_{\bm{\omega}}(.)$

### v4

@@ III Deep JSCC for data with unknown errors and distributions
In this section, the proposed robust and high-quality JSCC is further detailed. In Subsection III-A, WGAN and its inversion network are introduced to serve as the decoder and encoder. Subsection III-B then provides an fine-tuning encoder that is robust to errors. The fast and reliable SemCom approach for data of unknown distribution is implemented in Subsection III-C by latent space exploration for unknown distributions.
III-A Decoder and Encoder: GAN and GAN Inversion
The generators of GANs with slightly lower generation quality than DMs are still selected as the semantic decoder for the proposed LDM-enabled JSCC for its single-step data generation property. GAN is formulated based on zero-sum game between a discriminator $D_{\bm{\gamma}}(\cdot)$ and a generator $G_{\bm{\psi}}(\cdot)$ with the adversarial training objective given as follows:
$\min_{\bm{\psi}}\max_{\bm{\gamma}}$ $\mathbb{E}_{q(\bm{x})}\left[\log D_{\bm{\gamma}}(\bm{x})\right]+\mathbb{E}_{q_{\bm{\psi}(\bm{z})}}\left[\log\left(1-D_{\bm{\gamma}}(G_{\bm{\psi}}(\bm{z}))\right)\right],$ (4)
@@ III-A Decoder and Encoder: GAN and GAN Inversion
where $q(\bm{x})$ denotes the distribution of input data, $q_{\bm{\psi}}(\bm{z})$ is the prior distribution of latent vector $\bm{z}$ and $q_{\bm{\psi}}(\bm{z})=\mathcal{N}(\bm{0},\bm{I})$ in GANs. Furthermore, to overcome these challenges of training instability and collapse mode, WGAN [41] is established by replacing $\mathcal{D}_{KL}$ and Jensen-Shannon divergence $\mathcal{D}_{JS}$ with Wassertein distance $\mathcal{D}_{\mathcal{W}}$. Similarly, other variants of GAN that have been proposed for better perceptual reconstruction can also be utilized as the JSCC decoder. Among them, StyleGAN [51] and Diff-GAN distillated from DMs [52] have achieved impressive generation results, and Diff-GAN is even comparable to the DMs in some datasets.
@@
Compared to VAE, GANs are skilled in generating data with high-resolution. Nonetheless, the task of SemCom is to ensure that the signals received by receivers can accurately convey the meaning, while minimizing the bandwidth of SemCom. For this reason, JSCC requires an encoder to determine the latent bottlenecks of the transmitted data, also known as GAN inversion [53]. Commonly, the solution of GAN inversion is to utilize a neural network-based encoder to find the optimal latent vector $\bm{z}$ given the transmitted data $\bm{x}$.
Proposition 1.
Ignoring the channel’s cross entropy term of the latent space and taking into account the receiver reconstruction term and the transmitter encoding entropy, the VUB defined in Eq. (3) can be transformed into
@@
where the proof can be seen in Appendix -A.
Apparently, the term $\mathbb{E}_{q_{\bm{\psi}}(\bm{z})}\left[-\log p_{\bm{\psi}}(\bm{x})\right]$ in Eq. (5) can be replaced with the training objective of the generator $G_{\bm{\psi}}(\cdot)$ of WGAN, and term $\mathbb{E}_{q(\bm{x})}\left[\mathcal{D}_{KL}(q_{\bm{\phi}}(\bm{z}|\bm{x})\parallel p_{\bm{\psi}}(\bm{z}|\bm{x}))\right]$ indicates that the encoding latent vector $\bm{z}$ should be as consistent as possible with the input latent space of generator $G_{\bm{\psi}}(\cdot)$ under the same transmitted data $\bm{x}$, which can be addressed by training VAE. When DM can generate realistic $\bm{z}$ as much as possible, jointly or separately training VAE and WGAN is equivalent to minimizing the loss $\mathcal{L}_{JSCC}$. The output of VAE’s encoder can be represented as $q_{\bm{\phi}}(\bm{z}|\bm{x})\sim\mathcal{N}(\bm{\mu},\bm{\sigma}^{2})$, and $\bm{z}$ is reparameterized as $\bm{z}=\bm{\mu}+\bm{\sigma}\odot\bm{\epsilon}$, where $\bm{\epsilon}\sim\mathcal{N}(\bm{0},\bm{I})$ and $\odot$ denotes the element-wise product. Consequently, by combining the decoupled optimization objectives of WGAN and Eq. (5), the training process of deep CNN based VAE-WGAN can be found in [54] and Appendix -B.
III-B Robust Semantic Encoder
As discussed in Section II, VAE-WGAN based SemCom systems suffer from inevitable vulnerabilities. The adversarial attack methods explore the vulnerabilities of neural networks utilized for classification and regression tasks, with the goal of determining a sufficiently small and unnoticed error $\bm{\delta}$ to mislead the classification or regression results. The unified optimization objective of adversarial attacks is given by
@@ III-B Robust Semantic Encoder
s.t.: $\bm{f}(\bm{x}+\bm{\delta})=\mathcal{T},\quad\bm{L}\leq\bm{x}+\bm{\delta}\leq\bm{U},$
where $\bm{d}(\cdot,\cdot)$ is the distance function that measure the differences between two data points, $\bm{f}(\cdot)$ is the attacked nerual network, $\mathcal{T}$ denotes the output of the DL model, $\bm{L}$ and $\bm{U}$ represent the physical lower and upper bounds of input data $\bm{x}$, respectively. Specifically, in classification tasks, $\mathcal{T}$ is the class that is inconsistent with original category of $\bm{x}$, e.g., the hackers can exploit objective (6) to make their cyber attacks undetectable by network $\bm{f}(\cdot)$. In regression tasks, $\mathcal{T}$ represents the output data that is not the same as the original regression results, e.g., when transmitting a digital image in SemCom systems, decoder may reconstruct another type of digital image with semantic ambiguities.
Input: Dataset $q(\bm{x})$, learning rate $\eta_{1}$ and $\eta_{2}$, original encoder $E_{\bm{\phi}}(\cdot)$, generator $G_{\bm{\psi}}(\cdot)$,
Output: The updated robust encoder $E_{\bm{\phi}^{\prime}}(\cdot)$
@@ 10      Determine $\bm{\delta}$ by $\bm{\delta}\leftarrow\bm{\delta}^{k}$;
Return Robust GAN inversion $E_{\bm{\phi}^{\prime}}(\cdot)$
Algorithm 1 Training algorithm of robust GAN inversion $E_{\bm{\phi}^{\prime}}(\cdot)$
In order to address the challenges of semantic errors, SemCom systems should have a robust and enhanced encoder that can handle those outliers. The objective for the sufficiently small semantic error $\bm{\delta}$ constrained by $\varepsilon$ that leads to the maximum receiver reconstruction error is given by
where $\left\|\cdot\right\|_{p}$ denotes the p-norm. In this way, objective (7) simultaneously utilizes the vulnerabilities of both the encoder $E_{\bm{\phi}}(\cdot)$ and generator $G_{\bm{\psi}}(\cdot)$. When solving (7), its objective can be transformed into a standard convex optimization problem
where $\lambda$ is the penalty coefficient. The constrained convex optimization problem can be solved by using the projected gradient descent (PGD) [55] iterative optimization method to obtain semantic error $\bm{\delta}$. Consequently, the semantic error at the i-th iteration $\bm{\delta}^{i}$ is denoted by
where $P_{C}(\bm{\varsigma}^{i})$ represents the projection of $\bm{e}(\bm{\delta})$ on the set of constraints $C$, i.e., $\bm{\delta}^{i}=P_{C}(\bm{\varsigma}^{i}):={\arg\min}_{\bm{\delta}\in C}\frac{1}{2}\left\|\bm{\delta}-\bm{\varsigma}^{i}\right\|_{2}^{2}$.
Proposition 2.
Let $\bm{z}^{\prime\prime}$ be the erroneous latent vector encoded from data containing semantic errors $\bm{x}^{\prime}=\bm{x}+\bm{\delta}$, the robust VUB for semantic errors is defined as
$\mathbb{E}_{q}\left[-\log p_{\bm{\psi}}(\bm{x})\right]=\mathbb{E}_{q}\left[-\log\int p_{\bm{\psi}}(\bm{x},\bm{x}+\bm{\delta})d(\bm{x}+\bm{\delta})\right]$ (10)
$+\mathbb{E}_{q(\bm{z}^{\prime\prime})}\left[-\log p_{\bm{\psi}}(\bm{z}^{\prime\prime})\right]+\frac{\beta}{2}\mathbb{E}_{q(\bm{z},\bm{z}^{\prime\prime})}\bm{d}(\bm{z},\bm{z}^{\prime\prime})$
$-\mathbb{E}_{q(\bm{z},\bm{z}^{\prime\prime})}\left[-\log q(\bm{z},\bm{z}^{\prime\prime})\right],$
where $q(\bm{z},\bm{z}^{\prime\prime})$ denotes the joint distribution and the proof can be seen in Appendix -C.
Evidently, the first and second terms of Eq. (10) have been addressed in VAE-WGAN based JSCC, and the third term has also been optimized by sloving the semantic errors $\bm{\delta}$. For this reason, the training objective of robust encoder is
$\min_{\bm{\phi}^{\prime}}\quad\frac{\beta}{2}\mathbb{E}_{q(\bm{z},\bm{z}^{\prime\prime})}\bm{d}(\bm{z},\bm{z}^{\prime\prime})+\mathbb{E}_{q(\bm{z},\bm{z}^{\prime\prime})}\left[\log q(\bm{z},\bm{z}^{\prime\prime})\right],$ (11)
where $\bm{\phi}^{\prime}$ denotes the robust encoder parameters. In [56], the objective (11) is equivalent to minimizing the Wassertein distance between $\bm{z}$ and the incorrect $\bm{z}^{\prime\prime}$. Nevertheless, the ultimate goal of SemCom is to accurately reconstruct the transmitted data. Therefore, as the parameter $\bm{\psi}$ is fixed, the optimal robust encoder parameters $\bm{\phi}^{\prime}$ considering both encoder and decoder vulnerabilities is
III-C Out-of-Domain Latent Space
@@
DL-based SemCom will significantly degrade their performances when facing data types that are not included in the training dataset (out-of-domain). For the proposed JSCC approach, when the transmitter wants to send an out-of-domain data, the robust encoder $E_{\bm{\phi}^{\prime}}(\cdot)$ may encode an abnormal latent vector, and the decoder at the receiver will reconstruct data that is semantically different from the transmitted data. For this reason, when facing the data with unknown distributions, SemCom system should improve its generalization abilities and be able to quickly and dynamically adapt to search for the optimal out-of-domain latent space.
@@
To address this issue, a learning-based adaptor constructed by a lightweight single-layer neural network is utilized for out-of-domain latent space determination. Considering the characteristics of VAE-WGAN, as shown in Fig. 3, the adapter $g_{\bm{\omega}}(\cdot)$ parameterized by $\bm{\omega}$ is placed between the robust encoder and generator. Subsequently, when the transmitted data follows an unknown distribution, adaptor $g_{\bm{\omega}}(\cdot)$ can perform one-shot learning to transform the latent vector $\bm{z}$ encoded by $E_{\bm{\phi}^{\prime}}(\cdot)$ into
@@
where $\bm{b}$ denotes the bias of adaptor $g_{\bm{\omega}}(\cdot)$. In order to improve the data quality of reconstruction, inspired by the adversarial training startegy of WGAN, this paper considers another adaptor $d_{\bm{\nu}}(\cdot)$ composed of a lightweight fully connected (FC) layer for adversarial training with $g_{\bm{\omega}}(\cdot)$. As illustrated in Fig. 3, during online training, the FC layer of discriminator $D_{\bm{\gamma}}(\cdot)$ is replaced by $d_{\bm{\nu}}(\cdot)$, and the original discriminator after removing the FC layer is denoted by $d_{\bm{\gamma}}(\cdot)$. In this way, the online training process of $g_{\bm{\omega}}(\cdot)$ is similar to WGAN’s training approach as illustrated in Algorithm 2.
@@
Input: Data following unknown distribution $q(\bm{x}^{\prime\prime})$, learning rate $\eta$, gradient penalty coefficient $\lambda$, robust encoder $E_{\bm{\phi}^{\prime}}(\cdot)$, generator $G_{\bm{\psi}}(.)$, discriminator $D_{\bm{\gamma}}(\cdot)$ and the parameters of last layer (L-th layer) is denoted by $\bm{\gamma}^{L}$
Output: The online-updated adaptor $g_{\bm{\omega}}(\cdot)$